### v1

@@ Abstract
Synthetic Aperture Radar (SAR) object detection has gained significant attention recently due to its irreplaceable all-weather imaging capabilities. However, this research field suffers from both limited public datasets (mostly comprising <2K images with only mono-category objects) and inaccessible source code. To tackle these challenges, we establish a new benchmark dataset and an open-source method for large-scale SAR object detection. Our dataset, SARDet-100K, is a result of intense surveying, collecting, and standardizing 10 existing SAR detection datasets, providing a large-scale and diverse dataset for research purposes. To the best of our knowledge, SARDet-100K is the first COCO-level large-scale multi-class SAR object detection dataset ever created. With this high-quality dataset, we conducted comprehensive experiments and uncovered a crucial challenge in SAR object detection: the substantial disparities between the pretraining on RGB datasets and finetuning on SAR datasets in terms of both data domain and model structure. To bridge these gaps, we propose a novel Multi-Stage with Filter Augmentation (MSFA) pretraining framework that tackles the problems from the perspective of data input, domain transition, and model migration. The proposed MSFA method significantly enhances the performance of SAR object detection models while demonstrating exceptional generalizability and flexibility across diverse models. This work aims to pave the way for further advancements in SAR object detection. The dataset and code is available at https://github.com/zcablii/SARDet_100K.
Keywords:
Benchmark Computer Vision Object Detection Synthetic Aperture Radar Pretrain, Domain Transfer
1 Introdiction
Figure 1: (a) Advantages of SAR image: independent of weather conditions, sunlight and land cover. (b) Number of papers (in thousands) retrieved from Google Scholar using keywords “SAR Detection”.
Synthetic Aperture Radar (SAR) [57, 60] has emerged as a pivotal technology in remote sensing, providing numerous advantages over traditional optical sensors. Notably, SAR possesses the capability to acquire geographical images under any weather conditions, irrespective of factors such as sunlight, land cover, or certain types of camouflage, as demonstrated in Fig. 0(a). As a consequence of these advantages, SAR has found extensive applications in critical domains, including national defence [48], humanitarian relief [3, 68], camouflage detection [20], and geological exploration [51, 25].
@@ 1 Introdiction
To address this problem, we merge the most publicly available SAR detection datasets. This effort includes a comprehensive review of current public SAR detection resources, followed by the collection and standardization of these datasets into a uniform format, creating a unified large-scale multi-class dataset for SAR object detection, named SARDet-100k.
This dataset comprises approximately 117K images and 246k instances of objects across six distinct categories. To our knowledge, SARDet-100k is the first dataset of COCO-scale magnitude in this vital research area. It significantly contributes to overcoming the previously mentioned limitations by providing a rich resource for the development and evaluation of SAR object detection models. Moreover, the dataset and source code will be made publicly available.
Transferring gaps. Through our empirical research and detailed analysis, we have identified that a principal hurdle in SAR object detection is the significant domain gap and model gap encountered when transferring a backbone network pretrained on natural RGB datasets (e.g., ImageNet [18]), to a detection network on SAR detection datasets. The domain gap stems from the stark visual discrepancies between RGB and SAR imagery, whereas the model gap arises from the model differences between the pretrained backbone and the whole detection framework employed in the downstream task.
@@
The MSFA framework demonstrates remarkable efficacy in reducing the substantial domain and model gaps typically encountered between the pretraining and finetuning stages. Our proposed method is not only effective but also general and applicable across various modern deep neural networks.
@@
Figure 2: Visualization of handcrafted features on SAR images. (To facilitate visualization, the features are average pooled and represented as a single channel.)
@@ 2.1 SAR Images and Handcraft Features
SAR imaging often suffers from poor image quality due to multiplicative speckle noise and artifacts [57, 47]. To address this, many traditional handcrafted feature descriptors, such as HOG [16], Canny [5], GRE [30], Haar-like [62] and WST [45] feature descriptors are developed or adapted to extract more discernible features from SAR images. The following paragraphs will provide detailed discussions on each of these popular descriptors.
Histogram of Oriented Gradients [16] (HOG). HOG is a widely used local feature descriptor in image processing and computer vision. It captures and represents the image’s local structure and shape information by analyzing the distribution of gradient orientations. HOG is proven to be effective for SAR image classification and object detection tasks because it is invariant to random noise in SAR images. Early works employ HOG for SAR object recognition [56, 49], and recent studies demonstrate the effectiveness of HOG features for SAR image classification [82, 84], as well as the model pretraining [66] in modern neural networks.
Canny Edge Detector [5]. Canny is a widely used edge detection algorithm, which aims to identify significant edges in an image while minimizing noise and spurious responses. The algorithm utilizes Gaussian smoothing, pixel orientation gradient magnitude, and non-maximum suppression. Early research [29, 39] recognize the advantage of Canny for SAR image processing. Recent works also verify the effectiveness of the Canny feature in SAR image edge detection [69], interference detection [19], and image registration [85]. Other than Canny Edge Detector, Gradient by Ratio Edge (GRE) [30], a recently proposed edge detector, leverages SAR-HOG [17] and SAR-SIFT [56] to achieve effective edge detection in SAR images.
Haar-like [62] Feature Descriptor. Haar-like features are commonly used for face detection [46, 2], pedestrian detection [80], and other target detection tasks [33, 37]. They describe the characteristics of an image by utilizing predefined feature templates. Haar-like features can capture linear features, edge features, point features, and diagonal features. Early studies [54, 55] demonstrate the robustness of Haar-like features for SAR object detection. A recent approach called MSRIHL [1] integrates low-level Haar-like features into deep learning models for accurate SAR object detection, highlighting its potential effectiveness.
Wavelet Scattering Transform [45] (WST). The WST is a powerful signal processing technique that is widely used in image processing. It aims to extract robust and discriminative low-level and high-level features simultaneously. By capturing both high-frequency and low-frequency information, it provides a rich representation of both local and global image features. The hierarchical representation offered by the Wavelet Scattering Transform enables feature extraction at different scales and resolutions, which is particularly useful for capturing fine details of small objects in SAR images and robustly handling high-frequency noise. Vidal et al. [61] conduct a comprehensive investigation demonstrating the high potential performance of Wavelet Transform in SAR image denoising. Many recent works [92, 27, 50, 77] utilize Wavelet Transform or WST for robust feature extraction and integration with CNN networks for target recognition. These studies validate the feasibility of incorporating WST features into modern CNN models.
In recent years, the field of SAR image analysis has been predominantly influenced by deep learning approaches. Nevertheless, the potential of handcrafted features in this domain has largely been overlooked. While some previous studies attempt to integrate classic handcrafted features into modern neural networks for SAR image processing, they predominantly concentrated on tasks related to low-level [69, 85], classification [82, 84, 92, 27, 50, 77] and pretrain [30, 30]. The application of these features in domain adaptation and SAR object detection, however, remains largely untapped.
@@ 2.2 SAR Object Detection
Various popular deep learning-based object detection frameworks, including RetinaNet [34], FCOS [59], GFL [31], RCNN series [53, 4], YOLO series [52, 10], and DETR [6], demonstrate remarkable generalizability in the field of general object detection. Additionally, modern backbone networks such as ConvNext [43], VAN [23], LSKNet [32] and Swin Transformer [42] are designed to efficiently and effectively model visual features. However, SAR image object detection poses unique challenges due to factors such as small object size, speckle noise, and sparse information inherent in SAR images. Consequently, the direct transfer use of a general object detection model may not be suitable for SAR images. As a result, recent deep learning methods for SAR object detection primarily focus on network and module design to address these challenges. Approaches like MGCAN [9], MSSDNet [90], and SEFEPNet [78] enhance object features through multiscale feature fusion. Quad-FPN [81] combines four distinct feature pyramid networks for thorough multiscale feature interaction to alleviate noise interferences and multi-scale object feature misalignment. PADN [87] and EWFAN [65] employ attention mechanisms to enhance object features in the presence of SAR speckle noise. CenterNet++ [22], an extension of CenterNet [91], incorporates feature enhancement, multi-scale fusion, and head refinement modules to improve the detector’s robustness specifically for SAR images. Additionally, CRTransSar [74], built on the high-performance Swin transformer [42], leverages context representation learning to enhance object features.
While most existing works concentrate on mitigating SAR speckle noise interference through network structure improvements, few attempts to address the issue at the level of input data. Furthermore, many studies utilize ImageNet pretrained backbones as the initialization of the detection framework, overlooking the substantial domain gap between the pretrained nature scenes dataset and the finetuned SAR dataset, as well as the model gap between the backbone and the entire detection framework. Instead, we seek to address these unique challenges through a carefully designed pretraining strategy.
@@ 3 A New Benchmark Dataset for SAR Object Detection
In this section, we introduce a new large-scale benchmark for SAR object detection. We present a comprehensive SAR object detection dataset, which is a compilation of 10 existing datasets. This dataset will serve as the foundation for our experiments.
3.0.1 Current Status.
SAR images are typically captured by satellites, and there is a wealth of low-resolution SAR imagery available, often with a Ground Sample Distance (GSD) of 10m $\times$ 10m or larger. Platforms like Sentinel-1 [12] provide access to these images, which offer a macroscopic view of various geophysical places such as cities, mountains, rivers, and cultivated land. This makes them particularly advantageous for scene classification tasks. However, the inherent low resolution of these images constrains their capability to delineate fine details of smaller objects, such as ships, cars, and airplanes.
Conversely, high-resolution SAR images provide more detailed information but require significant hardware resources. Moreover, these high-definition images often encompass sensitive information, making them unsuitable for public release. Furthermore, acquiring high-resolution SAR datasets can be very expensive, posing significant challenges to their accessibility.
Numerous research teams frequently encounter budgetary limitations that restrict their capacity to obtain a large and diverse collection of high-resolution SAR dataset. These financial constraints not only limit the scope of geographical areas that can be covered but also affect the variety of data sources that can be accessed. Consequently, the datasets made available by these groups often lack diversity, particularly in aspects such as spectral bands, polarization, and resolution. From a researcher’s perspective, evaluating models on such small and homogeneous datasets can introduce bias and lead to unfair performance comparisons.
3.0.2 Dataset Collection.
To address the aforementioned challenges, we undertake a thorough survey of SAR object detection datasets. As a result, we collect a total of 10 publicly available high-quality datasets that are not only diverse but also have no conflicting object categories. These data are released by or collected from different countries and institutions, such as scientific research departments in China, space departments in Europe, and military departments in the United States. Detailed information on the collected datasets is shown in Table. 1.
Datasets Target Res. (m) Band Polarization Satellites
  AIR_SARShip S 1,3m C VV GF-3
HRSID S 0.5$\sim$3m C/X HH, HV, VH, VV S-1B,TerraSAR-X,TanDEMX
MSAR A, T, B, S $\leq$ 1m C HH, HV, VH, VV HISEA-1
SADD A 0.5$\sim$3m X HH TerraSAR-X
SAR-AIRcraft A 1m C Uni-polar GF-3
ShipDataset S 3$\sim$25m C HH, VV, VH, HV S-1,GF-3
SSDD S 1$\sim$15m C/X HH, VV, VH, HV S-1,RadarSat-2,TerraSAR-X
OGSOD B, H, T 3m C VV/VH GF-3
SIVED C 0.1,0.3m Ka,Ku,X VV/HH Airborne SAR synthetic slice
Table 2: Image and instance level statistics of SARDet-100K dataset. *: Origin datasets are cropped into 512 $\times$ 512 patches.
@@
3.0.3 Standardization.
To ensure consistency across the collected datasets, it is necessary to undertake standardization processes. This involves addressing variations in train-val-test splitting status, image resolutions, and annotation formats. The overview of SARDet-100K dataset process pipeline is illustrated in Fig. 2(a).
Furthermore, we convert all dataset annotations into the COCO annotation format [35]. This step ensures consistency and compatibility among the different datasets. Consequently, the merged dataset, SARDet-100K, is also standardized in the COCO format, which is readily compatible with popular open-source detection code frameworks, eliminating the need for additional manual data preprocessing.
Figure 3: (a) SARDet-100K dataset standardization process, encompassing set splitting, large image slicing, and label annotation format unification. (b) Percentage of instances for each category and average instance area (in pixels) in SARDet-100K.
3.0.4 SARDet-100K.
Table 2 presents the standardized sub-datasets of SARDet-100K along with their corresponding statistics, which includes information on both image-level and instance-level statistics. Fig. 2(b) provides an overview of the category-level statistics for the SARDet-100K dataset. The SARDet-100K dataset encompasses a total of 116,598 images, and 245,653 instances distributed across six categories: Aircraft, Ship, Car, Bridge, Tank, and Harbor. SARDet-100K dataset stands as the first large-scale SAR object detection dataset, comparable in size to the widely used COCO [35] dataset (118K images), which is the standard benchmark for general object detection. The scale and diversity of the SARDet-100K dataset provide researchers with robust training and evaluation for advancing SAR object detection algorithms and techniques, fostering the development of SOTA models in this domain.
@@ 4 Multi-Stage with Filter Augmentation Pretraining Framework
Several recent studies [27, 77, 74, 22] have demonstrated the effectiveness of mature handcrafted features and specialized network module designs in improving SAR object detection performance. However, most of these works rely on the default ImageNet pretraining approach, thus overlooking the significant domain gap between the pretrained nature scenes dataset and the finetuned SAR dataset. Additionally, they fail to address the model gap that exists between the backbone and the entire detection framework. To address these limitations, we propose a novel framework called the Multi-Stage with Filter Augmentation (MSFA) Pretraining Framework. Our framework tackles the challenges from the perspective of data input, domain transition, and model migration. MSFA comprises two core designs: the Filter Augmented Input and the Multi-Stage pretrain strategy.
@@
4.0.1 Filter Augmented Input.
@@
Where $T_{i}$ is a pre-defined transformation. The resulting Filter Augmented Input to the detection model, $Inp$, will be a concatenation of the original single channel grayscale SAR image $x$ and the generated filter augmented feature $M_{i}^{x}$ as:
@@
4.0.2 Multi-stage pretrain.
@@
Our proposed multi-stage pretrain strategy, alternatively, can be illustrated as in the Eq. (3) (5) and (6).
@@
Where an extra second stage pretrainin in Eq. (5) is added. We propose the utilization of a large-scale optical remote sensing dataset, $D_{RS}$, as a detection pretrain for domain transit. The dataset consists of optical modal imagery, which also shares similar object shapes, scales, and categories in downstream SAR datasets. This characteristic serves as a valuable bridge between the optical distribution of natural images in Imagenet and the object distribution in SAR remote sensing images. By leveraging such second stage pretrain, the domain gap is effectively minimized, as illustrated in Fig. 3(b).

### v2

@@ Abstract
Synthetic Aperture Radar (SAR) object detection has gained significant attention recently due to its irreplaceable all-weather imaging capabilities. However, this research field suffers from both limited public datasets (mostly comprising <2K images with only mono-category objects) and inaccessible source code. To tackle these challenges, we establish a new benchmark dataset and an open-source method for large-scale SAR object detection. Our dataset, SARDet-100K, is a result of intense surveying, collecting, and standardizing 10 existing SAR detection datasets, providing a large-scale and diverse dataset for research purposes. To the best of our knowledge, SARDet-100K is the first COCO-level large-scale multi-class SAR object detection dataset ever created. With this high-quality dataset, we conducted comprehensive experiments and uncovered a crucial challenge in SAR object detection: the substantial disparities between the pretraining on RGB datasets and finetuning on SAR datasets in terms of both data domain and model structure. To bridge these gaps, we propose a novel Multi-Stage with Filter Augmentation (MSFA) pretraining framework that tackles the problems from the perspective of data input, domain transition, and model migration. The proposed MSFA method significantly enhances the performance of SAR object detection models while demonstrating exellent generalizability and flexibility across diverse models. This work aims to pave the way for further advancements in SAR object detection. The dataset and code is available at https://github.com/zcablii/SARDet_100K.
1 Introduction
Synthetic Aperture Radar (SAR) [57; 60]is a pivotal technology in remote sensing, providing numerous advantages over traditional optical sensors. Notably, SAR possesses the capability to acquire geographical images under any weather conditions, irrespective of factors such as sunlight, land cover, or certain types of camouflage, as demonstrated in Fig. 1(a). As a consequence of these advantages, SAR has found extensive applications in critical domains, including national defence [48], humanitarian relief [3; 68], camouflage detection [19], and geological exploration [51; 24].
@@ 1 Introduction
Figure 1: (a) Advantages of SAR image: independent of weather conditions, sunlight and land cover. (b) Number of papers (thousands) retrieved from Google Scholar using keywords “SAR Detection”.
@@
To address this problem, we merge the most publicly available SAR detection datasets. This effort includes a comprehensive review of current public SAR detection resources, followed by the collection and standardization of these datasets into a uniform format, creating a unified large-scale multi-class dataset for SAR object detection, named SARDet-100k. This dataset comprises approximately 117k images and 246k instances of objects across six distinct categories. To our knowledge, SARDet-100k is the first dataset of COCO-scale magnitude in this research area. It significantly contributes to overcoming the previously mentioned limitations by providing a rich resource for the development and evaluation of SAR object detection models. Moreover, the dataset and source code will be made publicly available.
Transferring gaps. Through our empirical research and detailed analysis, we have identified that a principal hurdle in SAR object detection is the significant domain gap and model gap encountered when transferring a backbone network pretrained on natural RGB datasets (e.g., ImageNet [17]), to a detection network on SAR imagery. The domain gap stems from the stark visual discrepancies between RGB and SAR imagery, whereas the model gap arises from the model differences between the pretrained backbone and the whole detection framework employed in the downstream task.
@@
The MSFA framework demonstrates remarkable efficacy in reducing the substantial domain and model gaps typically encountered between the pretraining and finetuning stages. MSFA is not only effective but also general and applicable across various modern deep neural networks.
@@ 2.1 SAR Images and Handcraft Features
SAR imaging often suffers from poor image quality due to multiplicative speckle noise and artifacts [57; 47]. To mitigate this, many traditional handcrafted feature descriptors have been developed or adapted to extract more discernible features from SAR images. These include Histogram of Oriented Gradients [15] (HOG), Canny Edge Detector [5], Gradient by Ratio Edge (GRE) [29], Haar-like [62] Feature Descriptor and Wavelet Scattering Transform [45] (WST). Early works employed traditional algorithms, such as HOG for SAR object recognition [56; 49], Canny [28; 38] for edge detection. However, in recent years, the field of SAR image analysis has been largely dominated by deep learning approaches.
While recent studies have focused on tasks related to low-level processing [69; 86], classification [83; 85; 93; 26; 50; 78] and pretrain [29; 29], they have attempted to integrate classic handcrafted features into modern neural networks for robust SAR image feature extraction and refinement. In contrast, our work does not simply inject such handcrafted features into networks, but explores the benefits and potentials of handcrafted features in domain adaptation and SAR object detection under modern deep neural networks. This research area remains largely unexplored, and our work aims to bridge this gap.
@@ 2.2 SAR Object Detection
Various popular deep learning-based object detection frameworks, including RetinaNet [33], FCOS [59], GFL [30], RCNN series [53; 4], YOLO series [52; 10], and DETR [6], demonstrate remarkable generalizability in the field of general object detection. Additionally, modern backbone networks such as ConvNext [42], VAN [22], LSKNet [31] and Swin Transformer [41] are designed to efficiently and effectively model visual features. However, SAR image object detection poses unique challenges due to factors such as small object size, speckle noise, and sparse information inherent in SAR images. As a result, recent deep learning methods for SAR object detection primarily focus on network and module design to address these challenges. Approaches like MGCAN [9], MSSDNet [91], and SEFEPNet [79] enhance object features through multiscale feature fusion. Quad-FPN [82] combines four distinct feature pyramid networks for thorough multiscale feature interaction to alleviate noise interferences and multi-scale object feature misalignment. PADN [88] and EWFAN [65] employ attention mechanisms to enhance object features in the presence of SAR speckle noise. CenterNet++ [21], an extension of CenterNet [92], incorporates feature enhancement, multi-scale fusion, and head refinement modules to improve the detector’s robustness specifically for SAR images. Additionally, CRTransSar [74], built on the high-performance Swin transformer [41], leverages context representation learning to enhance object features.
While most existing works concentrate on mitigating SAR speckle noise interference through network structure improvements, few attempts to address the issue at the level of input data. Furthermore, most studies utilize ImageNet pretrained backbones as the initialization of the detection framework, overlooking the substantial domain gap between the pretrained nature scenes dataset and the finetuned SAR dataset, as well as the model gap between the backbone and the entire detection framework. Instead, we seek to address these unique challenges through a carefully designed pretraining strategy.
@@ 3 A New Benchmark Dataset for SAR Object Detection
3.1 Current Status
SAR images are typically captured by satellites, and there is a wealth of low-resolution SAR imagery available, often with a Ground Sample Distance (GSD) of 10m $\times$ 10m or larger. Platforms like Sentinel-1 [12] provide access to these images, which offer a macroscopic view of various geophysical places such as cities, mountains, rivers, and cultivated land. This makes them particularly advantageous for scene classification tasks. However, the inherent low resolution of these images constrains their capability to delineate fine details of smaller objects, such as ships, cars, and airplanes. Conversely, high-resolution SAR images provide more detailed information but require significant hardware resources. Moreover, these high-definition images often encompass sensitive information, making them unsuitable for public release. Furthermore, acquiring high-resolution SAR datasets can be very expensive, posing significant challenges to their accessibility.
Numerous research teams frequently encounter budgetary limitations that restrict their capacity to obtain a large and diverse collection of high-resolution SAR datasets. These financial constraints not only limit the scope of geographical areas that can be covered but also affect the variety of data sources that can be accessed. Consequently, the datasets made available by these groups often lack diversity, particularly in aspects such as spectral bands, polarization, and resolution. From a researcher’s perspective, evaluating models on such small and homogeneous datasets can introduce bias and lead to unfair performance comparisons.
Table 1: Image and instance level statistics of SARDet-100K dataset. *: Origin datasets are cropped into 512 $\times$ 512 patches. Ins: Instances, Img: Images.
@@ 3.1 Current Status
Datasets Target Res. (m) Band Polarization Satellites License
  AIR_SARShip [76] S 1,3m C VV GF-3 -
HRSID [71] S 0.5$\sim$3m C/X HH, HV, VH, VV S-1B,TerraSAR-X,TanDEMX GNU General Public
MSAR [75] A, T, B, S $\leq$ 1m C HH, HV, VH, VV HISEA-1 CC BY-NC 4.0
SADD [80] A 0.5$\sim$3m X HH TerraSAR-X -
SAR-AIRcraft [90] A 1m C Uni-polar GF-3 CC BY-NC 4.0
ShipDataset [67] S 3$\sim$25m C HH, VV, VH, HV S-1,GF-3 -
SSDD [84] S 1$\sim$15m C/X HH, VV, VH, HV S-1,RadarSat-2,TerraSAR-X Apache2.0
OGSOD [63] B, H, T 3m C VV/VH GF-3 -
SIVED [35] C 0.1,0.3m Ka,Ku,X VV/HH Airborne SAR synthetic slice -
3.2 SARDet-100K
To address the aforementioned challenges, we undertake a thorough survey of SAR object detection datasets. As a result, we carefully collect a total of 10 publicly available high-quality datasets that are not only diverse but also have no conflicting object categories. These data are released by or collected from different countries and institutions, such as scientific research departments in China, space departments in Europe, and military departments in the United States. Detailed information on the collected datasets is shown in Table. 2. To ensure consistency across the collected datasets, we invest considerable time and effort in rigorous dataset standardization. This involves addressing variations in train-val-test splitting status, image resolutions, and annotation formats. More details about data collection and standardization can be found in the Appendix.
Table 1 presents the standardized sub-datasets of SARDet-100K along with their corresponding statistics, which includes information on both image-level and instance-level statistics. The SARDet-100K dataset encompasses a total of 116,598 images, and 245,653 instances distributed across six categories: Aircraft, Ship, Car, Bridge, Tank, and Harbor. SARDet-100K dataset stands as the first large-scale SAR object detection dataset, comparable in size to the widely used COCO [34] dataset (118K images), which is the standard benchmark for general object detection. The scale and diversity of the SARDet-100K dataset effectively simulate real-world scenarios encountered in the application of SAR object detection models across multiple data sources. SARDet-100K provides researchers with robust training and evaluation for advancing SAR object detection algorithms and techniques, fostering the development of SOTA models in this domain.
@@ 4 Multi-Stage with Filter Augmentation Pretraining Framework
Several recent studies [26; 78; 74; 21] have demonstrated the effectiveness of mature handcrafted features and specialized network module designs in improving SAR object detection performance. However, most of these works rely on the default ImageNet pretraining approach, thus overlooking the significant domain gap between the pretrained nature scenes dataset and the finetuned SAR dataset. Additionally, they fail to address the model gap that exists between the backbone and the entire detection framework. To address these limitations, we propose a novel framework called the Multi-Stage with Filter Augmentation (MSFA) Pretraining Framework. Our framework tackles the challenges from the perspective of data input, domain transition, and model migration. MSFA comprises two core designs: the Filter Augmented Input and the Multi-Stage pretrain strategy.
@@
4.1 Filter Augmented Input
@@
Where $T_{i}$ is a pre-defined transformation. Drawing inspiration from the information residual design in ResNet [23], we construct the Filter Augmented Input to the detection model, $Inp$, by concatenating the original grayscale SAR image $x$ with the generated filter augmented feature $M_{i}^{x}$ as:
@@
4.1.1 Multi-stage pretrain
@@
Our proposed multi-stage pretrain strategy, alternatively, can be illustrated as in the Eq. (3) (5) (6).
@@
Where an extra second stage pretraining in Eq. (5) is added. We propose the utilization of a large-scale optical remote sensing dataset, $D_{RS}$, as a detection pretrain for domain transit. The dataset consists of optical modal imagery, which also shares similar object shapes, scales, and categories in downstream SAR datasets. This characteristic serves as a valuable bridge between the optical distribution of natural images in ImageNet and the object distribution in SAR remote sensing images. By leveraging such second stage pretrain, the domain gap is effectively minimized, as illustrated in Fig. 2(b).